4. Temperature

Assuming a filament in a 100W light bulb acts like a perfect blackbody, what is the temperature of the hottest portion of the filament if it has a surface area of6.3*10^-5m^2? The Stefan Boltzmann constant is 5.67*10^-8W/(m^2*K^2).
Answer #1
Power P = 100 Wsurface area A = 6.3 * 10 -5 m2Stefan Boltzmann constant σ = 5.67*10^ -8 W/(m^2*K^2)we know from stefan Law P = σ A T4from this temperature T = [ P / Aσ ] (1/ 4 )substitue values weget T value
answered by: Sumone